# Build An Influential Bot In Social Media Simulations With Large Language Models

### v1

@@ 5.1 Simulated Environment Initialisation
The initiation of our simulated environment begins with the creation of initial posts, for which we use the dataset released by Betz (2021). The dataset is crawled from debating platform kialo.com, comprises various posts structured around specific topics. For our simulation, we select posts related to the topic ’Gender’ and topic ’Drug’.
@@
Given the polarity distributions, we categorise the posts into distinct opinion groups, aiming for an even distribution across these categories. The categorisations is shown in table 4 and table 5.
@@ 5.2 Simulated Environment Analysis
In both scenarios, the distribution of the final polarity scores shows a slight skew towards higher values across the time steps. However, in the Narrow setting, the opinions form a more centralized distribution by step 50, indicating stronger clustering and a more definitive convergence of opinions, as compared to the Creative setting, which remains more dispersed throughout the simulation.
@@ 5.3 Reinforcement Learning Setting
By testing different setting, we found that limiting the action space and incorporating self-observation are key factors for stable RL result.
By restricting the action space, we reduce the complexity of the decision-making process, enabling the RL agent to focus on a more manageable set of actions. This helps prevent the model from exploring irrelevant or suboptimal actions that could introduce noise and instability into the learning process. Additionally, a smaller, well-defined action space allows for faster convergence, as the agent can more efficiently learn the consequences of its actions within the environment.
@@
For the reinforcement learning, we use Gender case and Drug case, set number of episodes 500, max of steps per episode as 5. Other parameters are shown in Appendix-Parameters.
5.4 Reinforcement Learning Result
This subsection presents the outcomes of our RL experiments conducted across the eight scenarios. Figure 5 illustrates the results for the Gender scenario and Figure 6 shows the results for the Drug scenario. In both figures, the colored lines represent the average reward of five repeated experiments, and the shaded areas indicate the variance in the results.
In all scenarios, we observe convergence in the reward learning curves, indicating that the RL algorithm successfully identified optimal solutions within the given settings. Both the Gender and Drug cases exhibit similar learning curve patterns under different conditions, highlighting consistent performance across varying contexts.
@@ 5.4 Reinforcement Learning Result
6 Conclusion
In this section, we reflect on our contribution, future work, ethical consideration, and limitations.
We have introduced a novel application of RL within a simulated discussion environment to model individual behavior in the context of social media. Our findings show that limiting the action space and incorporating self-observation are critical factors for generating stable results in dynamic environments. These insights demonstrate the adaptability of the RL model to diverse conditions and its effectiveness in identifying optimal strategies. Additionally, the RL results offer a deeper understanding of both individual and collective behavior on digital communication platforms.
Our future work would be directed towards enhancing the realism and variability of our simulated social media environment. This may involve incorporating more sophisticated interaction rules to better mirror the real-world social media dynamics. Additionally, we plan to discover how different character profiles of agents perform and interact within the network. This exploration aims to provide a deeper understanding of the varied behaviors and their impacts on the overall dynamics of social media interactions.
The potential application would be providing valuable insights for digital marketing strategists and policymakers in the dynamics of online opinion formation and influence.
However, we acknowledge that one potential negative social impacts could be the misuse of the simulated environment in manipulating public opinion. If used unethically, the technology may craft highly persuasive and targeted misinformation campaigns on social media platforms. Conversely, this technology can also be instrumental in detecting malicious activities in social networks.
@@ 7 Appendix A: GPT API Setting
Table 2 outlines the parameters set for the GPT API, distinguishing between ’creative’ case and ’narrow’ case configurations. Within these settings, higher temperature and top-p values are indicative of a setup geared towards fostering creativity. For tasks related to post generation and opinion elicitation, we employ the pre-trained GPT-3.5 model developed by OpenAI OpenAI (2022).
8 Appendix B: Reinforcement Learning Setting Parameters
Parameters for Reinforcement Learning training are listed in table 3.
learning rate 0.01
discount rate 0.99
epsilon 0.005
exploration rate 0.01
Table 3: RL Training Parameters
8.1 Appendix C: Opinion Example
@@
8.2 Appendix D: Opinion Category Setting

### v2

@@ 5.1 Simulated Environment Initialisation
The initiation of our simulated environment begins with the creation of initial posts, for which we use the dataset released by Betz (2022). The dataset is crawled from debating platform kialo.com, comprises various posts structured around specific topics. For our simulation, we select posts related to the topic ’Gender’ and topic ’Drug’. The dataset includes 486 posts under ’Gender’ topic and 660 posts under ’Drug’ topic, with each posts labeled with one type ’pro’ or ’con’. The length of each post is less than 70 words. For example, under ’Gender’ topic, a post such as ’The absence of gender constructions would enable people to enjoy a better overall quality of life.’ is labeled as type ’pro’.
@@
Based on the polarity distributions, we categorize the posts into distinct opinion groups while ensuring an even distribution across these categories. The specific range settings are detailed in Appendix D, and the categorization results are presented in Table 5 and Table 6.
@@ 5.2 Simulated Environment Analysis
Figure 5: Standard Deviation of Opinion on Narrow setting and Creative setting
@@
In both scenarios, the distribution of the final polarity scores shows a slight skew towards higher values across the time steps.
Figure 5 plots the standard deviation of opinion over time in both Narrow and Creative settings. The graph shows that in the Narrow setting, the opinions form a more centralized distribution by step 50, indicating stronger clustering and a more definitive convergence of opinions, as compared to the Creative setting, which remains more dispersed throughout the simulation.
@@ 5.3 Reinforcement Learning Setting
By testing different settings, we found that limiting the action space of the RL agent and incorporating self-observation are key factors for stable RL result. We found that actions when restricted by using only statements in the dictionary, and not the LLM, significantly improved convergence in training due to the narrow topic of our experiments (e.g., gender rights).
By restricting the action space, we reduce the complexity of the decision-making process (e.g., number of actions in a policy), enabling the RL agent to focus on a more manageable set of actions. This helps prevent the model from exploring either irrelevant or suboptimal actions within a policy that could introduce noise and instability into the learning process. Additionally, a smaller, well-defined action space allows for faster convergence, as the agent can more efficiently learn the consequences of its actions within the environment.
@@
We use Gender case and Drug case, set number of episodes 500, max of steps per episode as 5. Other parameters are shown in Appendix-Parameters.
5.4 Reinforcement Learning Results
This subsection presents the outcomes of our RL experiments conducted across the eight scenarios. Figure 6 illustrates the results for the Gender scenario and Figure 7 shows the results for the Drug scenario. In both figures, the colored lines represent the average reward (followers) of five repeated experiments, and the shaded areas indicate the variance in the results. The followers gained (reward) rises from 1 to 2 or 3 in different cases in a total agent population of 20. That is to say, we see a gain of 10% of the total available follower population.
Regarding a baseline, if there was no intelligence from RL and we just used initial conditions, it would be a statistical flat line with no followers gained over time.
In all scenarios, we observe convergence in the reward learning curves (followers), indicating that the RL algorithm successfully identified optimal solutions within the given settings. Both the Gender and Drug cases exhibit similar learning curve patterns under different conditions, highlighting consistent performance across varying contexts.
@@ 5.4 Reinforcement Learning Results
6 Conclusions and Future Work
Understanding the dynamics of public opinion evolution on online social platforms is crucial for understanding influence mechanisms and the provenance of information. Traditional influence analysis is typically divided into qualitative assessments of personal attributes (e.g., psychology of influence) and quantitative evaluations of influence power mechanisms (e.g., social network analysis).
Many of the analysis approaches assume how influence is spread (e.g., diffusion Ghaffarzadegan et al. (2024), opinion ODE Jin & Guo (2022), or stochastic bounded confidence models Siahkali et al. (2024)), but these are not topic-specific and lack dynamic adaptation at the individual scale. Our approach uses reinforcement learning (RL) to strategically decide what actions to take in conversations and an LLM to translate the actions into natural language. This enables us to better modulate our conversations to be natural and topic-specific, whilst maximizing the individual agent’s long-term influence upon peers. Whilst there is emerging work on integrating LLMs and RL agents Panagopoulos et al. (2024) and large-scale simulations of agent behaviour Park et al. (2024), we believe at time of writing this is an important contribution.
In terms of academic contribution, our work also enable researchers to ethically test new agents and create fine-scale experimentation on opinion dynamics for diverse topics.
Our current findings reveal the following: (1) RL can achieve a gain of 10% of the total available follower population, and (2) convergence can be better achieved by limiting the action space and incorporating self-observation. This demonstrates the simulation framework’s capacity to create agents that can adapt to complex and unpredictable social dynamics. The potential application would be providing valuable insights for digital marketing strategists and policymakers in the dynamics of online opinion formation and influence. However, we acknowledge that one potential negative social impacts could be the misuse of the simulated environment in manipulating public opinion. If used unethically, the technology may craft highly persuasive and targeted misinformation campaigns on social media platforms. Conversely, this technology can also be instrumental in detecting malicious activities in social networks.
@@ 6 Conclusions and Future Work
Our future work would be directed towards enhancing the realism and variability of our simulated social media environment. This may involve incorporating more sophisticated interaction rules to better mirror the real-world social media dynamics. To achieve this, first we plan to discover how different character profiles of agents perform and interact within the network. This exploration aims to provide a deeper understanding of the varied behaviors and their impacts on the overall dynamics of social media interactions. We can also improve the follower mechanism, where we can create more complex multi-topic link updates, where agreement on one topic, might lead to a following that is across all similar topics – reflecting closer to the way social connections form. This can be expanded to cover similar or shared geographies, hobbies and, lifestyles. Currently, we are limited by our real data sets on topics. We are in particular interested in any ability to model social acceptance of future emerging technologies (e.g., 6G), where there are no real data sets, but understanding social opinion is critical. We do note that there are emerging synthetic world simulations (e.g., academic Park et al. (2023) and commercial Project OdySSEy) to emulate a realistic life and social interactions.
7 Code & Data Availability Statement
The source code (Python) for running the simulations is published at https://github.com/AlminaJin/ InfluentialAgent. All data is available from public data bases.
8 Acknowledgments
The work is supported by: (1) "Networked Social Influence and Acceptance in a New Age of Crises" funded by USAF OFSR (FA8655-20-1-7031); and (2) "Communications Hub For Empowering Distributed ClouD Computing Applications And Research" funded by EPSRC/DSIT (EP/X040518/1) (EP/Y037421/1).
@@ 9 Appendix A: GPT API Setting
Table 4 outlines the parameters set for the GPT API, distinguishing between ’creative’ case and ’narrow’ case configurations. These parameters are used in LLMs to tune how much creativity they add to the natural language output. The role of LLM is only to map the RL opinion decision to a natural language format. Therefore, this creativity is only in the mapping stage, not in the decision stage. Most literature Park et al. (2022) use these parameters similarly to map the boundaries of the LLM performance. Hopefully this explains that this does not affect how opinion influence decisions are made, but only the text format in which it manifests. Within these settings, higher temperature and top-p values are indicative of a setup geared towards fostering creativity. For tasks related to post generation and opinion elicitation, we employ the pre-trained GPT-3.5 model developed by OpenAI OpenAI (2022).
9.1 Appendix B: Opinion Example
@@
9.2 Appendix C: Opinion Category Setting
Based on the polarity distributions, we categorize the posts into distinct opinion groups while ensuring an even distribution across these categories. The specific range settings and the categorization results are presented below.
@@ 9.2 Appendix C: Opinion Category Setting
9.3 Appendix D: Model Documentation
The source code (Python) for running the simulations is published at https://github.com/AlminaJin/InfluentialAgent.